OpenStudy (anonymous):

if a cirlce has an area of 36pi square feet find the circumference of the circle PLEASE HELP!

OpenStudy (anonymous):

12pi

OpenStudy (anonymous):

36pi = pi x r^2 r = 6 Circumference = 2pi x r = 12pi

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

\[36\pi^2 = \pi *r^2\], so if you simplfy "pi"s and then take square root of both sides, you get r=6.